OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x - 3y = -11 2x - 6y = -14 not good in english

OpenStudy (anonymous):

answers -1 0 1 2

OpenStudy (cj49):

5x-3y=-11 (multiply by 2) 2x-6y=-12 (multiply by 5)

jhonyy9 (jhonyy9):

5x-3y=-11 *2 2x-6y=-14 * - 5 10x -6y = -22 -10x +30y = 70 ------------------ 0 24y = 48 y= 48/24 y = 2 hope this is understandably easy